OpenStudy (emmynimmy):

Algebra Use the pair of functions to find f(g(x)) and g(f(x)). Simplify your answers. f(x) = x2 + 9, g(x) = Square root x + 6 f(g(x)) = g(f(x)) =

zepdrix (zepdrix):

Hey miss Emily c: \[\large\rm f=x^2+9\qquad\qquad\qquad g=\sqrt{x+6}\] So for this first one,\[\large\rm f(g)\]they want us to take all of g, and stick it place of the x in our f function,\[\large\rm f(g)=g^2+9\]See how I replaced the x with g? Ok but g is actually sqrt(x+6), so let's write our g that way,\[\large\rm f(g)=(\sqrt{x+6})^2+9\]

zepdrix (zepdrix):

simplify from that point.

OpenStudy (emmynimmy):

How do i simplify?

zepdrix (zepdrix):

square root and square are `opposite` or `inverse` operations of one another. So they "undo" one another.\[\large\rm (\sqrt{x+6})^2\quad=\quad x+6\]See how the square root and the square both disappear? Ya, they were able to undo one another. So that would be the first step in simplifying. Next combine any like-terms.

satellite73 (satellite73):

\[g(x)=x+5\] what is \(g(5)\)?

OpenStudy (emmynimmy):

x?

satellite73 (satellite73):

lol

satellite73 (satellite73):

\[g(x)=x+5\\ g(5)=5+5=?\]

OpenStudy (emmynimmy):

10? idk

OpenStudy (emmynimmy):

don't laugh at me lol

satellite73 (satellite73):

yes, 5+5=10

satellite73 (satellite73):

now \[f(x)=2x+3\]so \(f(10)=?\)

OpenStudy (emmynimmy):

20?

satellite73 (satellite73):

no, \(2\times 10=20\) you need \(2\times 10+3=20+3=?\)

OpenStudy (emmynimmy):

23

OpenStudy (emmynimmy):

im lost

satellite73 (satellite73):

whew that is your answer to the first one

OpenStudy (emmynimmy):

ahahaha "whew" see what i mean? its impossible.

satellite73 (satellite73):

no you aren't lost \[f(g(5))=f(1)=23\] so \(23\) is the answer

satellite73 (satellite73):

ok typo there \[f(g(5))=f(10)=23\]

satellite73 (satellite73):

get out the calculator for the next one ready?

OpenStudy (emmynimmy):

okay

satellite73 (satellite73):

\[C(s(h))=C(25h)=45\times 25h+560\] you have to compute \(45\times 25\)

OpenStudy (emmynimmy):

so 45 times 25?

OpenStudy (emmynimmy):

1125

satellite73 (satellite73):

i believe you put \(1125h+560\) in the first box

satellite73 (satellite73):

then for the second part, \(h=6\) use your calculator to compute \[1125\times 6+560\]

OpenStudy (emmynimmy):

7310

OpenStudy (emmynimmy):

ok next question

satellite73 (satellite73):

yes 7310

satellite73 (satellite73):

this a repeat?

OpenStudy (emmynimmy):

no question we haven't done yet

satellite73 (satellite73):

you put \(x=9\) get \[9^3-2-\frac{9}{9-10}\]

satellite73 (satellite73):

\[f(3)=3^2-3\times 3-4=?\]

OpenStudy (emmynimmy):

12

satellite73 (satellite73):

no actually \[9-9-4\neq 12\]

satellite73 (satellite73):

nice and slow for this one what is \(g(3)\)? if it is not clear, say so

OpenStudy (emmynimmy):

is that like 3 x 3

satellite73 (satellite73):

no, \(g\) is a set of ordered pairs, things that look like \((a,b)\) find the ordered pair in the list where the first number is 3, i.e. \((3,?)\) what is the second number?

OpenStudy (emmynimmy):

idk

satellite73 (satellite73):

oh dear, look with your eyes, there is nothing to compute here. where do you see a pair that starts with the number 3?

OpenStudy (emmynimmy):

Next:

OpenStudy (emmynimmy):